Given f(x)=x/3-x and g(x)=3x/1+x, find g(f(4)).
A. -4? B. -1? C. 1? D. 4?
Thank you :)

Answer by LinnW(1048)

f(x)=x/(3-x) and g(x)=3x/(1+x), find g(f(4))
We can first find f(4) = 4/(3-4) = 4/(-1) = -4.
So f(4) = -4
Now substitute -4 for x in 3x/(1+x)
3(-4)/(1+(-4))
-12/-3
4
So g(f(4)) = 4